OpenStudy (anonymous):

Find S8 for the series -2 + -10 + -50 + -250 +…

OpenStudy (tkhunny):

A third try? IF we are to assume that it is a Geometric Series, which the problem statement does NOT say, you can just add them up. Trivial a = a Boring a + ar = a(1+r) Now we're getting somewhere interesting! a + ar + ar^2 = a(1 + r + r^2) = \(a\cdot\dfrac{1 - r^{3}}{1-r}\) Generally a + ar + ar^2 + ... + r^{n-1} = a(1 + r + r^2 + ... + r^(n-1)) = \(a\cdot\dfrac{1 - r^{n}}{1-r}\) That's all it takes. Determine a = the 1st term r = the common ratio n = the desired sum rank and you are on your way.

OpenStudy (anonymous):

this is no help.. thnks anyway

OpenStudy (tkhunny):

Seriously? You're not going to do ANY work? What's the first term? You must be able to do that.

OpenStudy (tkhunny):

Come on!! Tell me the first term. Look up at the series and read it off.

OpenStudy (anonymous):

im kool... thnks again